# Equal contribution. * Corresponding author

### v1

@@ 4.1 Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset: one involving hand-object interaction and the other hand-virtual interaction, with a total of over 600 frames. To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 3, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [45, 27, 104]. The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 5. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subject ethnicities.
@@
Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation (defined by MANO [68]).
4.2 Benchmark Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cf. Tab. 1) in an automatic way following Sec. 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation, i.e., HTT [92] and A2J-Transformer [35], and reproduce them on our dataset for benchmark experiments. HTT [92] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [35] only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TheFormer-V baseline. We also exclude the additional action block used by HTT [92] for hand action prediction, focusing solely on hand pose estimation. We adjusted the anchor initialization phase of the A2J-Transformer [35] to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU for a fair comparison of their inference speed.
@@ 4.2 Benchmark Setup
4.3 Benchmark Results
Table 4: Benchmark results of TheFormer and state-of-the-art methods on thermal-based 3D hand pose estimation. Results are reported on the main testing set. $\uparrow$ denotes larger values are better, and vice versa. The fps displayed for sequence input indicates the number of sequences that models can process per second.
@@ 4.3 Benchmark Results
A2J-Transformer [35] Single 51.68 0.474 20.76 0.603 34
HTT [92] Single 49.09 0.489 20.69 0.599 211
TheFormer-S Single 48.25 0.510 22.60 0.565 120
HTT [92] Sequence 47.07 0.512 17.49 0.659 129
TheFormer-V Sequence 46.57 0.519 19.62 0.619 67
Comparison against State of the Arts. We compare the TheFormer against the state-of-the-art methods for thermal-based 3D hand pose estimation, as shown in Tab. 4. On the MEPE and AUC metrics, TheFormer-S outforms two competing methods under the single image-based setting, while TheFormer-V surpasses the counterpart HTT [92] given the same sequential images as input. Such an improvement stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cf. Sec. 3.3). A performance gap can be seen between single image-based and video-based settings for both HTT [92] and TheFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. We can also observe that TheFormer has slightly inferior results in the RA space relative to its opponents, leaving room for improvement in localizing individual hand joints once hand roots are correctly identified. This also suggests that our advancements are primarily attributed to our accurate localization of the overall hand especially the wrists. Thanks to our lightweight network design, TheFormer is highly efficient to run with fps of 120 and 67 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices.
Table 5: Comparison between different spectra. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
Method Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$
HTT - Sequence [92] RGB 43.30 0.542 15.43 0.697
Depth 41.62 0.559 17.66 0.654
NIR 41.57 0.562 16.41 0.679
Thermal 46.57 0.519 17.49 0.659
A2J-Transformer - Single [35] RGB 40.21 0.577 16.77 0.675
Depth 43.80 0.552 18.26 0.647
NIR 46.46 0.513 18.00 0.648
Thermal 51.68 0.474 20.76 0.603
Comparison between Spectrum. To compare the efficacy of different spectra (i.e., RGB, depth, NIR and thermal), we assess their performance by applying them as varying inputs to the HTT [92] and A2J-Transformer [35] method. As seen in Tab. 5, RGB spectrum yields favourable performance for both two methods as they provide rich multi-channel semantics, allowing for accurate hand pose reasoning. Specifically, A2J-Transformer [35] achieves comparable performance to HTT [92] with sequential input. This can be attributed to its usage of the larger ResNet50 [31] backbone that is pretrained with RGB images. However, this also increases the computation overhead, making A2J-Transformer [35] 6$\times$ slower than HTT [92] (cf. Tab. 4). Depth and NIR spectra show a marginal decrease in performance compared to RGB in well-illuminated conditions. However, the active illumination design of NIR sensors enables them to serve as supplements to RGB in dark conditions, as illustrated in Fig. 6. The thermal spectrum, operating independently of any active light source, still achieves plausible hand pose estimation accuracy, with a small discrepancy compared to other spectra. In addition to the above evaluation on the main testing set, we show more comparison of different spectra to highlight the value of thermal images under challenging conditions for 3D hand pose estimation.
4.4 Performance under Challenging Conditions
Figure 6: Examples of hand pose estimation results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on corresponding images.
@@ 4.4 Performance under Challenging Conditions
To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison of four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf. Tab. 2). Our automatic annotation pipeline (cf. Sec. 3.2) becomes infeasible since hands appear corrupted in either RGB or depth images. Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig. 6. As can be seen, RGB-based methods fail with gloves wearing (cf. Fig. 6 (a-b)) and in darkness (cf. Fig. 6 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrary to RGB techniques, as shown in Fig. 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
@@
NIR sensors are significantly disrupted by strong sunlight, affecting both NIR imaging and depth map creation, as shown in Fig. 6 (e-f). Conversely, thermal imaging is immune to sunlight and outdoor conditions. The temperature difference between hands and their surroundings in the thermal spectrum facilitates effortless identification of the hands, leaving the thermal-based estimator unaffected. The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications. Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
@@ 5 Limitation and Future Work
Despite the advancements, we acknowledge certain limitations here guiding our future work. First, the scale of our current dataset is small when compared to some large-scale datasets [18, 64, 52, 60] that contain millions of images (cf. Tab. 1). We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions to better support future investigations in this field. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other computer vision tasks relevant to human hands. In this regard, further efforts could be directed towards a) annotating the fine-grained hand action splits inside each sequence, b) leveraging the heat residual on objects after hand grasp to annotate the contact heatmap [7] and c) generating the ground truth mesh for both rigid [45] and articulated objects [18].
@@ 6 Conclusion
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TheFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.

### v3

@@ 4.1 Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 3.2) and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [43, 41, 37]. The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 5. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape.
Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
@@
4.2 Experiment Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cfTab. 1) automatically following Sec. 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT [15] and A2J-Transformer [46], and reproduced them on our dataset for experiments. HTT [15] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [46] only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT [15], focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer [46] to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
@@ 4.2 Experiment Setup
4.3 Thermal Image-based 3D Hand Pose Estimation
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods [15, 46] on the main testing set, as shown in Tab. 3. TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT [15] given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 3.3). A performance gap can be observed between single image-based and video-based settings for both HTT [15] and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT [15] for other spectra (cfTab. 4), proving its adaptability to different inputs.
@@ 4.3 Thermal Image-based 3D Hand Pose Estimation
(a) A2J-Transformer [46] Single 51.68 0.474 20.76 0.603 34
(b) HTT [15] Single 49.09 0.489 20.69 0.599 211
(c) TherFormer-S Single 44.64 0.539 18.34 0.643 136
(d) (c) w/o spatial transformer Single 48.79 0.491 20.15 0.609 174
(e) (c) w/o mask guidance Single 48.83 0.494 18.89 0.625 141
(f) HTT [15] Sequence 47.07 0.512 17.49 0.659 129
(g) TherFormer-V Sequence 43.36 0.549 17.36 0.661 52
Table 3: Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
Ablation Studies. We ablate our proposed mask-guided attention mechanism and the entire spatial transformer (cfSec. 3.3) to observe their impact. As can be seen in Tab. 3, the mask-guided attention mechanism notably contributes to TherFormer’s performance growth (row (c) vs. (e)), demonstrating its effectiveness in mitigating the effect of background clutter. Other components in the spatial transformer bring a large performance gain in the RA space (row (e) vs. (d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
4.4 Comparison between Spectrum
4.4.1 Qualitative Results under Challenging Conditions
@@
To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf. Tab. 1). Our automatic annotation pipeline (cf. Sec. 3.2) becomes infeasible since hands appear corrupted in either RGB or depth images. Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig. 6. We use the same model, i.e., HTT [15], trained on different spectra for a fair comparison. Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
RGB vs. Thermal. As can be seen, RGB-based methods fail with gloves wearing (cfFig. 6 (a-b)) and in darkness (cfFig. 6 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrary to RGB techniques, as shown in Fig. 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
@@
NIR, Depth vs. Thermal. NIR sensors are significantly disrupted by strong sunlight, affecting both NIR imaging and depth map creation, as shown in Fig. 6 (e-f). Conversely, thermal imaging is immune to sunlight and outdoor conditions. The temperature difference between hands and their surroundings in the thermal spectrum facilitates effortless identification of the hands, leaving the thermal-based estimator unaffected. The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications.
Figure 6: Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
4.4.2 Quantitative Results under Normal Conditions
To validate the versatility of thermal-based methods, we also evaluate their performance compared to other spectra (i.e., RGB, depth, NIR and thermal) under normal conditions for general-purpose use cases. We apply different spectral images as input to multiple baselines methods and report the results in Tab. 4. Not surprisingly, depth image-based methods yield the best performance on average, as they directly utilize the depth information to provide a detailed 3D structure of the hand. The NIR spectrum shows a marginal decrease in performance compared to depth but outperforms both RGB and thermal spectra. This can be attributed to NIR camera’s active sensing ability, which leads to more consistent and reliable imaging regardless of the variability of external illuminations (vs. RGB) and temperature (vs. thermal). However, active NIR sensors are prone to interference from external NIR sources like sun glare (cfFig. 6). Thermal-based methods, despite using single-channel heat capture, achieves comparable performance to those using RGB images that contain rich color and textures. Such results suggest that thermal images can not only serve as supplements in challenging cases but also can be a viable alternative to other spectra in normal conditions for hand pose estimation.
HTT (Sequence) TherFormer-V TherFormer-S Average
Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$
RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.80 0.538
Depth 41.62 0.559 39.70 0.581 39.84 0.579 40.39 0.573
NIR 41.57 0.562 40.79 0.575 40.98 0.575 41.11 0.571
Thermal 47.07 0.512 43.36 0.549 44.64 0.539 45.06 0.533
Table 4: Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
@@ 5 Limitation and Future Work
Despite the advancements, we acknowledge certain limitations here guiding our future work. First, the scale of our current dataset is relatively small when compared to some datasets [29, 40, 39, 38]. We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts will be annotating the fine-grained hand action splits and contact heatmap for hand grasp [42]. Lastly, we currently rely on qualitative analysis for challenging conditions when our automatic annotation method becomes unreliable. In the future, we will leverage unaffected spectral images in individual cases to manually annotate the 3D hand pose.
@@ 6 Conclusion
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.
Acknowledgments and Disclosure of Funding
This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1), and partially supported by Cisco Research.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We elaborate our contributions as a list in the introduction section and briefly explain them in the abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Sec. 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] There is no societal impact of this work.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have reviewed the ethics review guidelines and make sure we obey all rules for this research.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] This work does not contribute to the theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] This work does not contribute to the theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [No] We will release our code upon the acceptance of this paper.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We describe them in our main paper and supplementary document.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] In line with the state-of-the-art 3D hand pose estimation works, we do not report error bars for our experiment results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We include the type of GPUs in the main paper.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We use the offical code of HTT [15] and A2J-transformer [46] to reproduce their results on our dataset, which have been cited.
    2. (b)
      Did you mention the license of the assets? [Yes] We mention the license of the offical codes of HTT and A2J-Transformer in the supplementary document.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include the datasheet and the download URL of our dataset in the supplemental document.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We claim in the footnote that participant consent forms were signed before the collection.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in the supplementary document.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We do not give written instructions to our participants.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] There is no potential participant risks.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our participants are volunteer.